OpenStudy (anonymous):

Identify the zeros of the function: y=(x-13)(x+14)

OpenStudy (angela210793):

x=13 or x=-14 cause x-13=0 when x=13 and x+14=0 when x=-14 and (x-13)*(x+14)=0 when one of those is 0 or both